OpenStudy (anonymous):

Write the equation of the line that is perp. to the line y= 3/4X +3 and passes through the point (3,2)

OpenStudy (anonymous):

Perp means your slope is the negative reciprocal so in the equation y=mx+b your m (which is slope) is -4/3

OpenStudy (anonymous):

I'm going to let kayc4 finish cause I have to go good luck though!

OpenStudy (anonymous):

when it is perpendicular, the slope will be the opposite inverse of the original so it will be -4/3. plug -4/3 and the given point into point slop formula to get y-2 = -4/3(x-2). then distribute simplify and set equal to y

OpenStudy (anonymous):

so I came out with y= -4/3x + 14/3

OpenStudy (anonymous):

Thank you so much ! the explanation helped so much ! makes a lot more sense now!